# The Exchangeability Assumption for Permutation Tests of Multiple Regression Models

### v1

@@ 2.3 Why permutation test?
    There do exist settings where mathematical and permutation approaches produce different inferential results, even within the linear model framework (e.g., see Section 4). Knowledge of permutation-based approaches (plus computational skills) allows inference in such areas where linear model assumptions don’t hold, and permutation tests may be straightforward to implement in situations where typical parametric tests may exist but be extremely complicated. As long as permuted draws can be made from the appropriate null world, permutation tests can be used to make inferences without parametric assumptions.
3 Full example analysis
A full simulation study comparing permutation methods is both redundant (of the literature, see Anderson and Legendre (1999) and Winkler et al. (2014)) and outside of the goals of our paper. However, we provide an analysis of a single real dataset to demonstrate the methods and compare the resulting p-values.
The data come from the National Bridge Inventory (NBI) [^1] https://www.fhwa.dot.gov/bridge/nbi.cfm. Subject to the National Bridge Inspection Standards, each state is responsible for preparing and maintaining an inventory of all bridges. NBI provides data for New York in 2023222https://www.fhwa.dot.gov/bridge/nbi/2023/delimited/NY23.txt with a complete coding guide [^3] https://www.fhwa.dot.gov/bridge/mtguide.pdf. In what follows, we examine a small random sample of bridges in New York that go over highways and are made of either steel or concrete. Our interest is in modeling improvement cost (proposed cost of improvement) as a linear function of average daily traffic volume ($X_{1}$) and the material ($X_{2}$) of the bridge (either concrete or steel). Figure 1 provides a visual representation of the linear relationship between cost and volume, with a shift due to the type of material.
In order to assess whether or not the type of material (steel or concrete) is a significant predictor variable in the multiple linear regression model of improvement cost on volume and type of material, we run five different hypothesis tests (all with the same null hypothesis that the coefficient on type of material is zero, i.e., $H_{0}\!\!:\beta_{2\cdot 1}=0$).
$E[\texttt{cost}]=\beta_{0\cdot 1,2}+\beta_{1\cdot 2}\texttt{volume}+\beta_{2\cdot 1}\texttt{material}$
The five models are OLS and the four permutation methods described above in Section 2. Table 2 displays the p-values calculated by each of the five different hypothesis tests. Some important aspects to note from the analysis are:
  • The average daily traffic volume is significant. (The p-value is not shown, but the relationship is seen in Figure 1).
  • The two explanatory variables (average daily traffic volume and type of material) are potentially associated, as seen in Figure 2.
  • The data are not symmetrically distributed about the regression lines, violating the normality condition of ordinary least squares.
  • All of the models require that the observations are independent.
Despite multiple potentially important model condition violations (e.g., normality of residuals and homoscedasticity of residuals), all five p-values (see Table 2) are on the same order of magnitude and produce the same inferential conclusions about the coefficient on type of material ($X_{2}$).
Figure 1: For a small sample of bridges in New York state, the cost for improvement seems to be a function of both the average daily traffic volume of the bridge as well as the bridge material (steel or concrete).
Table 2: A comparison of the resulting p-values obtained from applying ordinary least squares and each of the permutation methods to data on bridges in New York state with 2000 permutations. The inference seeks to understand the effect of construction material (steel or concrete) on cost of repairs in a regression model adjusted by traffic volume.
permutation type p-value
treatment variable (Draper & Stoneman) 0.14
response variable (Manly) 0.16
reduced model residuals (Freedman & Lane) 0.13
full model residuals (ter Braak) 0.13
OLS 0.15
Figure 2: The average daily traffic volume seems potentially associated with the type of material used in the bridge, with concrete bridges having a higher median traffic volume.
4 Extensions to account for clustered data
The bridges example demonstrates how permutation testing can be used to undertake inference in standard multiple regression model settings. As we have seen above, ideas of exchangeability are reasonably straightforward in the MLR model, and we can walk through the details to understand how the exchangeability conditions impact the appropriate conclusions (see Table LABEL:tbl-perms).
However, permutation methods are particularly attractive since they can be adapted to more complicated settings where exchangeability might not be as straightforward. In particular, one of the conditions for inference in multiple regression is that the observations (or more accurately, the residuals) are independent of one other. For the bridges setting, we might speculate that the observations are not independent. For example, consider bridges in a particular area (e.g., a city) that reflect underlying features, which might systematically affect traffic flow and/or cost (e.g., availability of resources or contracts with particular construction companies). We might imagine that two bridges in the same city would share a common construction market and patterns of use that are associated with the cost.
Violations of independence show up in many types of studies. For example, consider a situation where, for convenience or design purposes, families might be sampled rather than individuals. Principled and flexible methods have been developed for multiple linear regression models in the setting where the observations are clustered (see for example Laird and Ware (1982)). Random intercept models, where each observational unit (e.g., family) is assumed to have an underlying (unobserved) random level, are commonly used to account for clustering. However, such mixed effects models are both complicated (mixed effects models are not typically seen in undergraduate curricula) and may require sufficient sample size to assume asymptotic normality of regression coefficients (Maas and Hox 2005).
4.1 Cluster permuting
We consider an example where we are interested in predicting an outcome ($Y$) based on a quantitative predictor ($X_{1}$) and a dichotomous treatment ($X_{2}$) where some of the families have two individuals within the sample. An adaption of the permutation tests that accounts for clustering within observational unit is a computational alternative to the mathematical approach, which maintains appropriate Type I error rate (Winkler et al. 2014, 2015). Here we explore the clustered permutation idea with three different data scenarios, representing independent observations and two different types of clustering that violate the independent observation requirement.
4.1.1 Permuting independent observations
Figure 3 displays a simple example where the observations are independent and there is no clustering structure. There are eleven subjects, each from a different family; six of the subjects are exposed and five of the subjects are control samples.
In the example, we permute $X_{2}$, akin to the method of N. Draper and Stoneman (1966). We note that because we are permuting, there are always five exposed observations and six control observations, leading to (${11\choose 6}$ = 462 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). We fit a multiple linear regression on the quantitative predictor and the permuted treatment and use the resulting null distribution of the t-statistics to assess the significance of the treatment variable.
Figure 3: Each color and row represents a family structure. There are 11 individuals, none of them are in the same family. Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure.
4.1.2 Permuting homogeneous clustered observations
In the homogeneous clustered (e.g., same value of the treatment within a family) data scenario, there are multiple observations of the same type within the observational unit (see Figure 4). There are a total of eight families. Five of the families have a single observation, and three of the families have two individuals.
With homogeneous clustered observations, the permutation scheme must align with the family structure. Here, each family has only one treatment (either exposure or control), so we permute the treatments with the constraint that three of the single-person families and one of the two-person families are control. For the homogeneous clustered example and scenario, there are (${5\choose 3}\cdot{3\choose 1}$ = 30 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). After each permutation, we fit a multiple linear regression on the quantitative predictor and the permuted treatment. The resulting null distribution of the t-statistics is used to assess the significance of the treatment variable.
Figure 4: Each color and row represents a family structure. There are 5 individual “families” and three families with two members each. In the original structure treatment happens by family (e.g., air pollution near / not near the family home). Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure. The families are permuted together so as to keep intact the original scenario of treatment within family.
4.1.3 Permuting heterogeneous clustered observations
With heterogeneous clustered observations, the permutation scheme must align with the family structure. Here, the treatment on individual families will be permuted as if the observations were independent (i.e., a permutation with the constraint that two individuals are exposed and three are control). The permutation of observations in the two-member families happens within the family. That is, with equal probability, the treatment levels are either kept the same or swapped, within each family.
Figure 5: Each color and row represents a family structure. There are 5 individual “families” and three families with two members each. In the original structure, for the families of size two, exactly one member of the family is exposed, and one member of the family is in the control group. Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure. For the families of size two, the permutation happens within the family so as to keep intact the original structure of both treatments across the family.
4.2 Cluster simulations
We undertook a limited series of simulation studies to explore the behavior of permutation tests when the design structure determined a violation of the independence condition. These simulations exploited the fact that ignoring clustering for the homogeneous clusters yields standard errors that are too small and that doing so for the heterogeneous cluster design, where family members serve as their own control, yields standard errors that are too large (Cannon et al. 2001).
The heart of our simulations is well described by Winkler et al. (2015) who focuses on
For each model/scenario combination, 2000 simulations were generated with n = 20 subjects per group (more subjects than in Figure 3, Figure 4, and Figure 5). For the clustering scenarios, half of the subjects were singletons and the others shared a family member within the group (homogeneous) or across groups (heterogeneous). Within each simulation the null distribution was created using 2000 permutations.
Table 3: Type I error rates for different scenarios (2000 simulations, each with 2000 permutations for each simulation). Models include multiple linear regression (lm), multiple linear mixed effects (lme), permutation not accounting for clustering (naïve), and permutation accounting for clustering (correct). Forms of clustering are independent (yellow), homogeneous (white), and heterogeneous (blue). The dichotomous treatment variable was permuted (Draper and Stoneman, 1996).
Model/Scenario Type I error rate (95% CI) Results
lm independent 0.052 (0.043-0.063) matches desired alpha level
correct permutation independent 0.054 (0.045-0.065) matches desired alpha level
lm homogeneous 0.08 (0.068-0.092) anticonservative
lme homogeneous 0.048 (0.04-0.059) matches desired alpha level
naïve permutation homogeneous 0.081 (0.069-0.094) anticonservative
correct permutation homogeneous 0.055 (0.045-0.066) matches desired alpha level
lm heterogeneous 0.025 (0.019-0.033) conservative
lme heterogeneous 0.044 (0.036-0.054) matches desired alpha level
naïve permutation heterogeneous 0.03 (0.023-0.038) conservative
correct permutation heterogeneous 0.054 (0.045-0.065) matches desired alpha level
Table LABEL:tbl-simresults displays the results from the simulation study. As expected, the models that do not appropriately account for clustering are anticonservative (“lm homogeneous”, “naïve permutation homogeneous”) or overly conservative (“lm heterogeneous” and “naïve permutation heterogeneous”); the models are given in boldface. Methods that appropriately account for the clustering (e.g., “lme” or “correct permutation”) yielded the desired Type I error rate. When there are cluster dependencies, the naïve method is not advised.
@@ 5 Discussion and pedagogical recommendations
Our work presents some of the existing literature on permutation tests in the multiple linear regression setting. While it is not immediately obvious how to best permute an MLR model, it turns out that, generally, the different methods perform similarly with respect to Type I errors (and power). However, the difficult and extremely powerful concept of exchangeability can be accessed through the MLR setting in the classroom. After digging into the exchangeability ideas for the MLR case, we have found that students are able to apply a permutation approach to more complicated data settings (which are often encountered in real applications), like clustered observations. This guidance is consistent with the advice of Chance et al. (2024), who hoped “to help instructors understand the differences in the simulation strategies to better inform their own decisions of how to adapt a simulation approach for their classes and to better respond to student questions that may arise”.
@@
We close with some practical guidance about how to bring exchangeability and permutation tests into the classroom:
  • Teach permutation tests for MLR. Talk to your students about the choice they will have to make (permute $Y$? $X_{2}$? residuals?) and that all modeling contains choices. Choices are not usually objective or unbiased, so an ability to defend their choice is what gives them power as a statistician.
  • Create a simulation program (Rmd/qmd file or Shiny App) where students can change the error structure and discover the desired Type I error rates are not achieved only when the errors are particularly egregious.
  • Describe the difference between creating a null sampling distribution (i.e., making sure that the null hypothesis is true) and establishing that exchangeability holds (i.e., making sure that the Type I error rates will be accurate).
  • Present some or all of the MLR methods described here (with more in Winkler et al. (2014)) and have students fill out a blank Table LABEL:tbl-perms.
@@
  • Come up with scenarios (or use our cluster scenarios!) where the students can figure out the correct permutation schemes. See an example from JH’s class using a stratified two-sample test. https://st47s.com/Math154/Notes/permschp.html#macnell-teaching-evaluations-stratified-two-sample-t-test
@@
The ideas we have put forward can be taught at a variety of levels. Whether to permute $Y$ or $X_{2}$ can be taught as early as introductory statistics where students have done some hypothesis testing and some multivariate modeling. Graduate students can think carefully about how to prove that conditions of exchangeability are met. But the group of students who might most benefit from the ideas we’ve presented include upper level undergraduates and early graduate students who are focused on understanding which models are best in which settings and how to differentiate those settings. We believe that the discussions which will come from our presentation will be interesting and fruitful for broad statistical learning.

### v2

@@ 3.1 Why permutation tests?
    There do exist settings where mathematical (i.e., based on the Central Limit Theorem) and permutation approaches produce different inferential results, even within the linear model framework. Knowledge of permutation-based approaches (plus computational skills) allows inference in such areas where linear model assumptions don’t hold, and permutation tests may be straightforward to implement in situations where typical parametric tests may exist but be extremely complicated. As long as permuted draws can be made from the appropriate null world, permutation tests can be used to make inferences without parametric assumptions.
3.2 Set-up
The population model of interest is
$E[Y|X_{1},X_{2}]=\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2},$
where $X_{1}$ is the variable of interest in predicting $Y$, and $X_{2}$ is a nuisance variable. The $\cdot$ (dot) notation indicates the other predictor variables included in the model.
The model estimated from the original dataset is given by
$\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
An equivalent framing to the original data model is given for the permuted data. For example, if $Y$ is permuted to get $Y^{*}$, then the model estimated from the permuted dataset is given by
$\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
where $b^{*}_{1\cdot 2}$ is the sample coefficient on $X_{1}$ given that $X_{2}$ is in the model, and $b^{*}_{2\cdot 1}$ is the sample coefficient on $X_{2}$ given that $X_{1}$ is in the model, while it is $Y^{*}$ being regressed on $X_{1}$ and $X_{2}.$
All standard errors of the coefficients ($SE(b)$ and $SE(b^{*})$) are calculated using the normal ordinary least squares (OLS) formula used in standard linear regression software. Additionally, it is worth pointing out that underlying all of the methods (including OLS), the observations are assumed to be independent of one another.
3.3 Comparison of permutation methods
Unlike SLR, with MLR, there is no obvious choice of how to permute. In what follows, we discuss four different permutation methods and their exchangeability conditions. The goal of permuting is to create a null sampling distribution of the statistic of interest, here $b_{1\cdot 2},$ so that we may infer the variability of the statistic. That is, a distribution of the statistics under the setting where the outcome $Y$ and the predictor of interest $X_{1}$ are not linearly related. But also, the null distribution needs to be created under the exchangeability condition—that the permuting only leads to a change in the relationship between $Y$ and $X_{1}$ without changing any of the other variable relationships in the linear model. All four of the permuting schemes set the null hypothesis to $H_{0}\!\!:\beta_{1\cdot 2}=0.$
Statistics that are independent of all unknown parameters are called pivotal statistics. For permutation tests in the MLR setting, there are two reasons that we use t-statistics, which are pivotal or asymptotically pivotal (Winkler et al. 2014; ter Braak 1992). First, pivotal statistics allow for comparison across all methods, even the method that permutes the residuals of the full model and does not include $\beta_{1\cdot 2}=0$ in the test statistic. Second, the advantages of pivotal statistics are well-established (Hall and Titterington 1989; Hall and Wilson 1991; Westfall and Young 1993).
Using pivotal statistics, we dive into different permutation choices for the MLR model in order to help students develop a deeper understanding of the connection between exchangeability and how the permutation is implemented. Communicating the different permutation structures is an ideal way to explore the ideas of exchangeability. That is, even though the methods below are quite similar with respect to Type I errors and power, the discussion of the methods allows students study exchangeability and permutations more generally. We include extended details of the permutation schemes in Appendix 1. Here we briefly describe the four permutation methods with specific thought to the exchangeability violations for each one.
3.3.1 Permute $Y$
At first glance, it might seem like permuting the outcome variable would be a good way to break the relationship between $Y$ and $X_{1}$ (Manly 1986, 1997) (see complete algorithm in Section 1). Indeed, permuting $Y$ will break the relationship between $Y$ and $X_{1}$, which will force the null hypothesis to be true (which is what we want for testing). However, permuting $Y$ will also simultaneously break the relationship between $Y$ and $X_{2},$ which may not be acceptable if we need to preserve the relationship to mirror the original data structure. Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $Y$. Note that when $Y$ is permuted, the original relationship between $X_{1}$ and $X_{2}$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{2}$ and $Y$, however, is problematic in terms of exchangeability. If we permute $Y,$ then exchangeability is met only if $Y$ and $X_{2}$ are uncorrelated in the original dataset.
Table 2: Different permutation schemes, variable relationships that are broken, and variable relationships that are preserved. Any relationships that are broken and not null violate exchangeability. Note that all methods require the observations to be independent of one another.
Permutation Broken Relationships Preserved Relationships Permutation distribution
Permute $Y$ $X_{1}$ $Y$ $X_{1}$ $X_{2}$ $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$
Manly (1986), Manly (1997) $X_{2}$ $Y$ See Eq (1)
Permute $X_{1}$ $X_{1}$ $X_{2}$ $X_{2}$ $Y$ $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$
N. Draper and Stoneman (1966) $X_{1}$ $Y$ See Eq (2)
Permute reduced model residuals $X_{1}$ $Y$ (if $X_{1}$ $X_{2}$ are uncorrelated) $X_{1}$ $X_{2}$ $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$
Freedman and Lane (1983) $X_{2}$ $Y$ See Eq (3)
Permute full model residuals None $X_{1}$ $X_{2}$ $t^{*}=\frac{b_{1\cdot 2}^{*}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}$
ter Braak (1990) $X_{1}$ $Y$ See Eq (4)
ter Braak (1992) $X_{2}$ $Y$
3.3.2 Permute $X_{1}$
In order to maintain the relationship between $Y$ and $X_{2}$ (while still interested in the relationship between $Y$ and $X_{1}$), we might consider permuting $X_{1}$ instead of $Y$ (N. Draper and Stoneman 1966) (see complete algorithm in Section 1). Indeed, the permutation distribution created from permuting $X_{1}$ will force the null hypothesis to be true. However, permuting $X_{1}$ has the side effect that the relationship between $X_{1}$ and $X_{2}$ will be broken in the permuted data. If the data come from, for example, a randomized clinical trial (where $X_{1}$ is the treatment variable), then $X_{1}$ and $X_{2}$ will be independent in the original dataset, and permuting of $X_{1}$ will not violate the exchangeability condition. If $X_{1}$ and $X_{2}$ are correlated in the original dataset, as seen in our AP course and socio-economic status example, then permuting $X_{1}$ violates the exchangeability condition. Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $X_{1}$. Note that when $X_{1}$ is permuted, the original relationship between $X_{1}$ and $Y$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{1}$ and $X_{2}$, however, is problematic under general conditions.
3.3.3 Permute reduced model residuals
In order to address some of the concerns of the first two MLR permutation methods, Freedman and Lane (1983) propose a permutation scheme based on residuals (see complete algorithm in Section 1). The permutation preserves the relationship between $X_{1}$ $X_{2}$ as well as the relationship between $X_{2}$ $Y$. However, in order for the relationship between $X_{1}$ $Y$ to be broken (i.e., to obtain a null sampling distribution for the test of $H_{0}:\beta_{1\cdot 2}=0),$ $X_{1}$ and $X_{2}$ must not be associated. Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the reduced model residuals. Note that when permuting the reduced model residuals, the original relationships between both $X_{1}$ $X_{2}$ and additionally between $X_{2}$ $Y$ are preserved (which is what we want in terms of exchangeability). The null hypothesis is true only if the relationship between $X_{1}$ and $Y$ is broken, and that happens only when $X_{1}$ and $X_{2}$ are uncorrelated. Appendix Section 2 sketches the dependence of $\rho(Y^{*},X_{1})$ (the correlation between $Y^{*}$ and $X_{1}$) on $\rho(X_{1},X_{2})$ (the correlation between $X_{1}$ and $X_{2}).$
3.3.4 Permute full model residuals
As an extension to Freedman and Lane (1983), ter Braak’s (1990, 1992) permutation method permutes the residuals from the full model (see complete algorithm in Section 1). Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the full model residuals. Note that permuting the residuals under the full model allows all of the exchangeability conditions to hold. The new model does not force the null hypothesis to be true, which is why the test statistic measures the deviation of the permuted coefficients (on $X_{1}$), $b_{1\cdot 2}$, to the original data model coefficient (on $X_{1}$), $b^{*}_{1\cdot 2}$, instead of comparing to zero.
@@ 4 Discussion and pedagogical recommendations
In the specific context of the MLR model, there are violations of exchangeability (e.g., clustering) which affect performance and violations of exchangeability (e.g., the correlation between $X_{1}$ and $X_{2}$) which do not affect performance. Our deep dive into permutation tests for MLR is meant to communicate ideas of exchangeability mathematically and pedagogically. We acknowledge that we do not have a smoking gun example which shows which of the standard MLR permutation methods is “best” (indeed, they are roughly equivalent procedures), but we find their introduction to students as a helpful structure to explore exchangeability in a meaningful way. However, an advantage of permutation tests is that they give accurate inferences even in small samples (Anderson and Legendre 1999).
Our work presents some of the existing literature on permutation tests in the MLR setting. While it is not immediately obvious how to best permute an MLR model, it turns out that, generally, the different methods perform similarly with respect to Type I errors (and power). However, the difficult and extremely powerful concept of exchangeability can be accessed through the MLR setting in the classroom. After digging into the exchangeability ideas for the MLR case, students are able to apply a permutation approach to more complicated data settings (which are often encountered in real applications), like clustered observations. This guidance is consistent with the advice of Chance et al. (2024), who hoped “to help instructors understand the differences in the simulation strategies to better inform their own decisions of how to adapt a simulation approach for their classes and to better respond to student questions that may arise”.
@@
We close with some practical guidance about how to bring exchangeability and permutation tests into the classroom. The suggestions are organized by what type of classroom might be most appropriate.
introductory undergraduate
  • Teacher creates a simulation program (Rmd/qmd file or Shiny App) where students can change the error structure within an MLR analysis and discover the desired Type I error rates are not achieved only in the case when the errors are particularly egregious.
intermediate / advanced undergraduate
  • Teach permutation tests for MLR. Talk to your students about the choice they will have to make (permute $Y$? $X_{1}$? residuals?) and that all modeling contains choices. Choices are not usually objective or unbiased, so an ability to defend their choice is what gives them power as a statistician.
  • Describe the difference between creating a null sampling distribution (i.e., making sure that the null hypothesis is true) and establishing that exchangeability holds (i.e., making sure that the Type I error rates will be accurate). Have students describe which aspect of the permutation does which job.
  • Come up with scenarios (or use our cluster scenarios!) where the students can figure out the correct permutation schemes. See an example from JH’s class that uses permutations to address a research question using a stratified two-sample test. https://st47s.com/Math154/Notes/permschp.html#macnell-teaching-evaluations-stratified-two-sample-t-test
  • Carry out a variety of permutation tests using a multiple linear regression model from an example dataset (e.g., the bridges data in the supplementary materials). Have students check the implementations or provide some of the implementations and have them carry out the others.
advanced undergraduate
@@
  • Ask students to read the paper by Helwig (2019a) and explore the implications for the MLR setting via the nptest package in R (Helwig 2023). Have them explore how and why robust permutation tests (DiCiccio and Romano 2017) use alternative test statistics.
advanced undergraduate / beginning graduate
  • Present some or all of the MLR methods described here (with more in Winkler et al. (2014)) and have students fill out a blank Table LABEL:tbl-perms. Ask students to report which relationships are broken and which are preserved.
@@
We suggest some ideas at particular levels, but many of the pedagogical ideas can be adjusted to be effective at a variety of levels. Whether to permute $Y$ or $X_{1}$ can be taught as early as introductory statistics, where students have done some hypothesis testing and some multivariate modeling. Graduate students can think carefully about how to prove that conditions of exchangeability are met. The group of students who might most benefit from the ideas we’ve presented include upper level undergraduates and early graduate students who are focused on understanding which models are best in which settings and how to differentiate those settings.